# Robust Counterfactual Explanations under Model Multiplicity Using Multi-Objective Optimization

### v1

@@ Abstract
In recent years, explainability in machine learning has gained importance. In this context, counterfactual explanation (CE), which is an explanation method that uses examples, has attracted attention. However, it has been pointed out that CE is not robust when there are multiple machine-learning models. These problems are important when using machine learning to make safe decisions. In this paper, we propose robust CEs that introduce a new viewpoint—Pareto improvement—and a method that uses multi-objective optimization to generate it. To evaluate the proposed method, we conducted experiments using both simulated and actual data. The results demonstrate that the proposed method is robust and useful. We believe that this research will contribute to a wide range of research areas, such as explainability in machine learning, decision-making, and action planning based on machine learning.
@@ 1 Introduction
In this paper, we propose a robust CE for model multiplicity by introducing a new viewpoint, Pareto improvement, and multi-objective optimization. In addition, we devised a validation index and verified its robustness through the validation of simulated data and its usefulness through the validation of real data. The proposed method has four key features. The first is the use of the concept of Pareto improvement to address the problem of model multiplicity; the second is that it can be used in practice to select safe solutions based on CE for risky problems; the third is that it can be applied not only to class classification but also to continuous target variables, such as in regression, and is sufficiently versatile to allow the free inclusion of constraint conditions; and fourth, by extracting a variety of CEs, it is possible to select a solution according to the user’s preference.
@@ 3 Experiment
In Experiment 1, the robustness of the CE extracted by the proposed method was verified using simulated data; in Experiment 2, it was applied to real data, and its usefulness was discussed.
@@ 3.1 Experiment 1: Simulation Data
  • Method 2: Multiple CEs are generated by changing the initial values based on a stacking model (multiple regression) using the predictions of the above models. The generation method is the same as described above.
@@
The parameters were set as $\lambda=2$ and $C=3$. For the calculation of val, since $y_{t}=\infty$, this study simply used the predicted values of $y$, where higher values indicated better evaluations. Moreover, as val and dissim or plaus tended to increase together, the ratios of their average values, ave val/ave dissim and ave val/ave plaus, were calculated to compare their balance. For Method 3, the average of multiple ave val values was used as ave val. The top three lowest values for dissim and plaus, and the top three highest values for FIR, ave val/ave dissim and ave val/ave plaus, are highlighted in bold.
@@
Additionally, we aimed to infer the impact of $\lambda$ and $C$. The value of $\lambda$ affected the balance between val and dissim, where a decrease in one typically led to an increase in the other. As the value of $C$ increased, the range of possible CEs expanded, leading to an increase in FIR across all methods.
@@ 3.2 Experiment 2: Real Data
The proposed method is applied to real-world data, where the true model is unknown. This allowed us to investigate the impacts of specific variables, evaluate their usefulness, and examine their potential applications in practical scenarios.
@@
Figure 2: Average of averages of CEs, excluding baseline data
@@
Figure 3: Pareto front
@@ 4 Discussion
In this study, we investigated a robust CE for machine learning, particularly for the problem of model multiplicity, which has become an issue in recent years. Specifically, we introduced the concept of Pareto improvement for a robust CE against model multiplicity and proposed the extraction of a robust CE using MOO. In addition, we propose a robustness index.
@@
This study also shows that the concept of Pareto improvement and multi-objective optimization can be applied to a wide range of areas, such as explainability in machine learning, decision-making, and action planning based on machine learning.
@@
Another potential avenue for future research is to incorporate causality to enhance the model’s robustness [10]. Causality is related to the invariant structure of the underlying data, which can ultimately contribute to robustness.

### v3

@@ Abstract
In recent years, explainability in machine learning has gained importance. In this context, counterfactual explanation (CE), which is an explanation method that uses examples, has attracted attention. However, it has been pointed out that CE is not robust when there are multiple machine-learning models with similar accuracy. These problems are important when using machine learning to make safe decisions. In this paper, we propose robust CEs that introduce a new viewpoint—Pareto improvement—and a method that uses multi-objective optimization to generate it. To evaluate the proposed method, we conducted experiments using both simulated and real data.The results demonstrate that the proposed method is both robust and practical. This study highlights the potential of ensuring robustness in decision-making by applying the concept of social welfare. We believe that this research can serve as a valuable foundation for various fields, including explainability in machine learning, decision-making, and action planning based on machine learning.
@@ 1 Introduction
In this paper, we propose robust CEs for model multiplicity by introducing a new viewpoint, Pareto improvement, and multi-objective optimization. In addition, we devised a validation index and verified its robustness through the validation of simulated data and its practicality through the validation of real data. The proposed method has four key features. The first is the use of the concept of Pareto improvement to address the problem of model multiplicity; the second is that it can be used in practice to select safe solutions based on CE for risky problems; the third is that it can be applied not only to class classification but also to continuous target variables, such as in regression, and is sufficiently versatile to allow the free inclusion of constraint conditions; and fourth, by extracting a variety of CEs, it is possible to select a solution according to the user’s preference.
@@ 3 Experiment
In Experiment 1, the robustness of the CE extracted by the proposed method was verified using simulated data; in Experiment 2, it was applied to real data, and its practicality was discussed.
@@ 3.1 Experiment 1: Simulation Data
  • Method 2: Multiple CEs are generated by changing the initial values based on a stacking model (multiple regression) using the predictions of the above models. The generation method is the same as described above.This method can also be interpreted as a solution approach for multi-objective optimization using weighted linear summation of objective functions, provided that all coefficients are positive.
@@
The parameters were set as $C=3$ and $\lambda=2$. For the calculation of val, since $y_{t}=\infty$, this study simply used the predicted values of $y$, where higher values indicated better evaluations. Moreover, as val and dissim or plaus tended to increase together, the ratios of their average values, ave val/ave dissim and ave val/ave plaus, were calculated to compare their balance. For Method 3, the average of multiple ave val values was used as ave val. The top three lowest values for dissim and plaus, and the top three highest values for FIR, ave val/ave dissim and ave val/ave plaus, are highlighted in bold.
@@
Additionally, we aimed to infer the impact of $C$ and $\lambda$.The value of $C$ increased, the range of possible CEs expanded, leading to an increase in FIR across all methods. As the value of $\lambda$ affected the balance between val and dissim, where a decrease in one typically led to an increase in the other. Regarding the ave val/ave dissim , we cannot assert a single, definitive trend because the functions may increase or decrease simultaneously, or there may be a trade-off relationship between them.
@@ 3.2 Experiment 2: Real Data
The proposed method is applied to real-world data, where the true model is unknown. This allowed us to investigate the impacts of specific variables, evaluate their practicality, and examine their potential applications in practical scenarios.
@@
(vertical axis: average; error bars: standard deviation)
Figure 2: Average of averages of CEs, excluding base data
(vertical axis: average; error bars: standard deviation)
@@
Figure 3: Pareto front(red: base data; blue: CEs)
@@
(vertical axis: average; error bars: standard deviation)
@@ 4 Discussion
In this study, we investigated robust CEs for machine learning, particularly for the problem of model multiplicity, which has become an issue in recent years. Specifically, we introduced the concept of Pareto improvement for robust CEs against model multiplicity and proposed the extraction of robust CEs using MOO. In addition, we propose a robustness index.
@@
This study highlights the potential to ensure robust decision-making by applying Pareto improvement—a concept from welfare economics—together with multi-objective optimization. We believe this research can serve as a valuable foundation for various fields, including explainability in machine learning, decision-making, and action planning based on machine learning.
@@
Another potential avenue for future research is to incorporate causality to enhance the model’s robustness [10]. Causality pertains to the invariant structure of the underlying data, which can ultimately contribute to improved robustness. We would also like to examine the applicability of concepts such as utilitarianism and Rawlsianism to the problems addressed in this study.